OpenStudy (anonymous):

Find the exact value of tan(cos^-1(-4/5))

OpenStudy (anonymous):

need help please

OpenStudy (jdoe0001):

\(\bf tan\left(cos^{-1}\left(-\frac{4}{5}\right)\right)\\ \color{blue}{cos^{-1}\left(-\frac{4}{5}\right) = \theta \implies cos(\theta) = -\cfrac{4}{5} \implies \cfrac{a}{c}\\ c^2= a^2+b^2 \implies \sqrt{c^2-a^2} = b}\\ tan\left(cos^{-1}\left(-\frac{4}{5}\right)\right) \implies tan(\theta) \implies \cfrac{b}{a}\)

OpenStudy (jdoe0001):

I was thinking on what quadrant the angle is, but ..... we only know that a = -4 but that could be the 2nd and 3rd quadrants so I gather there are 2 answers to the tangent function

OpenStudy (anonymous):

ahaa

OpenStudy (anonymous):

ahaa

OpenStudy (anonymous):

ammm okey so is the first answer -3/4 ?

OpenStudy (jdoe0001):

more like \(\bf \pm \cfrac{3}{4}\)

OpenStudy (anonymous):

hhmmmm and r u sure about it?

OpenStudy (jdoe0001):

you see the pythagorean theorem doesn't tell you if a root is either positive or negative, that will depend on the context in this case the context is, we have an angle whose cosine is negative any angle on the 2nd Quadrant will have a negative cosine any angle on the 3rd Quadrant will have a negative cosine also previously you knew where the angle was, so you knew what sign "b" will be

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

so you see, the cosine for those 2 angles is equally -4/5

OpenStudy (anonymous):

i did it the right answer is -3/4 it's correct (((:

OpenStudy (anonymous):

Thank you very much jdoe i learned a lot from you ^^

OpenStudy (jdoe0001):

yw